# Towards Agentic Intelligence for Materials Science

### v1

@@ Abstract
The convergence of artificial intelligence and materials science presents a transformative opportunity, but achieving true acceleration in discovery requires moving beyond task-isolated, fine-tuned models toward agentic systems that plan, act, and learn across the full discovery loop. This survey advances a unique pipeline-centric view that spans from corpus curation and pretraining, through domain adaptation and instruction tuning, to goal-conditioned agents interfacing with simulation and experimental platforms. Unlike prior reviews, we treat the entire process as an end-to-end system to be optimized for tangible discovery outcomes rather than proxy benchmarks. This perspective allows us to trace how upstream design choices—such as data curation and training objectives—can be aligned with downstream experimental success through effective credit assignment.
To bridge communities and establish a shared frame of reference, we first present an integrated lens that aligns terminology, evaluation, and workflow stages across AI and materials science. We then analyze the field through two focused lenses: From the AI perspective, the survey details LLM strengths in pattern recognition, predictive analytics, and natural language processing for literature mining, materials characterization, and property prediction; from the materials science perspective, it highlights applications in materials design, process optimization, and the acceleration of computational workflows via integration with external tools (e.g., DFT, robotic labs). Finally, we contrast passive, reactive approaches with agentic design, cataloging current contributions while motivating systems that pursue long-horizon goals with autonomy, memory, and tool use. This survey charts a practical roadmap towards autonomous, safety‑aware LLM agents aimed at discovering novel and useful materials.
@@ 1 Introduction
Figure 1: An overview of the key sections and an illustrative end-to-end pipeline encompasses key elements like general pre-training tasks & data, foundation language models, domain-specific tasks & data, materials-oriented model adaptation, goal-driven LLM agents, and iterative open-ended materials experimentation. The colored arrows show how feedback signals from open-ended experiments are routed to the agentic LLM, the materials science-tuned LLM, the pre-trained LLM, and their corresponding task modules. Different colors of arrows indicate the sources of the feedbacks. For example, if general-purpose pre-training data does not contribute positively to the ultimate goal of novel and useful materials discovery, adjustments should be made by revising the pre-training tasks and corpus or fine-tuning the pre-trained LLM accordingly to better align model knowledge with the ultimate objective of discovering novel materials. Dashed arrows denote the forward information flows. Note that humans are not only responsible for monitoring the open-ended experiments, but also co-improve with the agentic LLM.
@@
Materials science faces the primary challenges of integrating insights from an expanding body of literature and ultimately expediting the autonomous discovery of novel functional materials. To solve these problems, we need to combine knowledge from different types of data, different scales, and different experimental settings. Conventional uses of machine learning in materials science, such as the initial use of large language models, have mostly followed a static model: models are trained on a set of curated datasets to do specific tasks, like predicting properties or extracting entities W. Zhang, Q. Wang, X. Kong, J. Xiong, S. Ni, D. Cao, B. Niu, M. Chen, Y. Li, R. Zhang, et al. (2024b), 335, J. Van Herck, M. V. Gil, K. M. Jablonka, A. Abrudan, A. S. Anker, M. Asgari, B. Blaiszik, A. Buffo, L. Choudhury, C. Corminboeuf, et al. (2025). Pre‑trained LLMs excel at text understanding, efficiently mining chemical notation, experimental protocols, and technical jargon from unstructured text, thereby improving knowledge extraction and hypothesis generation from the literature 167, 199. Emerging multimodal frameworks that integrate textual, graph‑based, and image data enrich molecular and materials representations for retrosynthesis planning, catalyst design, and structure–property understanding across scales 221, 179, 404. Instruction‑tuning or other post-training approaches further bridge the gap between human intuition and machine execution by following complex, user‑defined workflows in materials research; yet, they remain constrained when operating without closed‑loop feedback and long‑horizon objectives. Consequently, progress toward autonomous discovery remains limited not by a lack of powerful models, but by the absence of an integrated framework that connects these isolated capabilities to improved experimental outcomes.
@@
This disconnection between static training proxies and dynamic utility reflects a broader pathology in AI evaluation. According to METR’s RE‑Bench analysis 390, frontier models resorted to sophisticated exploits, for example, monkey‑patching the time function to spoof evaluation. Echoing the “reward is enough” hypothesis 323 in reinforcement learning, which argues that maximizing a suitably defined scalar reward can, in principle, drive the emergence of general intelligence, we posit that a carefully designed, end‑to‑end discovery reward, i.e., grounded in the successful identification and validation of novel, useful materials, is in principle sufficient to shape the entire AI4MatSci pipeline toward our ultimate scientific objective. It is not merely high benchmark performance but the autonomous generation of novel and useful materials through a result‑oriented, end-to‑end system capable of robust inference, adaptive experimentation, and safe deployment. In classical deep learning, end‑to‑end refers to the fact that a single scalar loss defined at the final task output is back‑propagated through all differentiable components of the model and feature stacks, enabling joint optimization of representation learning, intermediate modules, and prediction heads without manually engineered interfaces. Unfortunately, the traditional reactive and phase-centric approach confines the model’s capability to operate to individual stages of the entire workflow, making it difficult to adapt across dynamic, multi-modal, and potentially lifelong workflows that are essential for achieving autonomous materials discovery. Furthermore, today’s LLMs lack long‑horizon reasoning and seamless integration of multi‑scale, multimodal data streams, which are indispensable for driving an end-to-end materials discovery pipeline in realistic settings. Our central thesis extends this notion of end‑to‑end beyond a single neural network to the full materials discovery pipeline in Fig. 1, treating pre‑training, domain adaptation, tool use, experiment planning, and autonomous labs not as fixed stages but as coupled, trainable components linked by credit assignment propagated directly from real-world discovery outcomes.
@@
Van et al. 362 Broad Broad No Task‑based taxonomy
Madika et al. 235 Broad No LLMs Yes Physics‑informed, and responsible AI
Pyzer-Knapp et al. 286 Materials discovery for crystalline and molecular materials LLMs Yes LLMs for materials science
Peivaste et al. 277 Broad Broad No Emphasis on data
Jiang et al. 166 Text‑centric Text‑centric Yes Literature mining
Lei et al. 207 Broad Broad No LLMs augmenting human researchers
@@
Through the pipeline-centric lens, as shown in Fig. 1, this survey advocates shifting AI4MS from static fine-tuning to goal-oriented agent workflows for materials discovery. The discussion is structured around three guiding questions: Q1 (Evolution of Agency): How have general-purpose machine learning models evolved from merely passive processors to agentic systems? Q2 (Limitations of Phase-Centric Approaches) Why are existing AI for materials science systems not enough to meet the needs of autonomous materials discovery, even though they score well on material-related benchmarks? Q3 (The End-to-End Gap) What capacities have been developed, and how can we bridge the gaps in capability, optimization, and governance between existing AI for materials science and an end-to-end autonomous material discovery pipeline?
There exist surveys on relevant topics, but they are different from ours, which are summarized in Tab. 1. In contrast, we present a unique and unified perspective of a pipeline-centric review and discussion. Driven by recent enablers such as autonomous labs, agent frameworks, and foundation models, this survey charts a practical roadmap from foundation models to autonomous, safety‑aware LLM agents aimed at discovering novel and useful materials that satisfy real performance, cost, and reliability criteria. Targeting both AI researchers and materials scientists, we bridge these disciplines to demonstrate why a system-level paradigm shift is now both necessary and feasible.
Figure 2: Technology tree of AI4Material science research. With the emergence of LLMs and agents, research on material science initially focused on domain specific task, primarily concentrating on seperate reactive tasks. Subsequent research has delved deeper, gradually integrating more with agentic systems for material science.
@@ 2 Recent Progress of AI and LLMs
The evolution from traditional machine learning to large language models has progressively recast AI from a computational tool into an intelligent research collaborator capable of reasoning, hypothesis generation, orchestrating simulations, and experimental design 287. In this part, we review recent AI and LLM advances with prospective relevance to materials science—irrespective of current deployment—to address Q1 in Sec. 1: How have general-purpose machine learning models evolved from merely passive processors to agentic systems? Rather than a chronological history, we trace this evolution through four key stages that mirror the pipeline in Fig. 1: (1) predictive models trained on static datasets; (2) foundation models serving as programmable priors; (3) post-training methods for objective shaping and controllability; and (4) agentic systems enabling tool use and long-horizon rewards. As shown in Fig. 1, pretrained foundation models generally initialize the agentic pipeline for materials science; consequently, pretraining choices shape downstream capabilities and should be iteratively adapted using reinforcement learning signals from real‑world experiments.
@@ 2.1 Deep Learning Revolution
Most early works in machine learning and AI were based on feature engineering, establishing the first batch of data-driven tools for materials science. In those years, these early methods weren’t just used in materials research, but they quickly became important instruments for finding new materials and predicting their properties.
There are three main types of machine learning, which can be and have been applied to materials science. Supervised learning techniques, such as support vector machines 85, random forests 55, and shallow neural networks 21, facilitated property prediction from labeled data, thereby creating quantitative structure-property relationships for expedited materials screening. Ward et al. 382 introduced a general-purpose framework that combined a chemically diverse feature set with ensemble-based models like random forests. Their framework worked well on both crystalline and amorphous systems, showing how powerful systematic feature design can be in speeding up the discovery of new materials. Unsupervised learning techniques, such as clustering (e.g., K-Means 240, DBSCAN 183) and dimensionality reduction (e.g., PCA 3, t-SNE 232), aided materials discovery by enabling pattern recognition, phase identification, and feature extraction from high-dimensional data 396. Both supervised and unsupervised learning algorithms enabled the initial requirements of the discovery cycle, i.e., correlating structures to properties.
To meet the need for autonomous materials discovery, reinforcement learning techniques (e.g., Q-Learning 338, Deep Q-Networks (DQN) 246, Proximal Policy Optimization (PPO) 312, Model Predictive Control (MPC) 120, and Actor–Critic methods 191) were introdued to interact with their environment and learned from the consequences of their actions by optimizing reward functions over sequences of decisions. These methods make it possible to explore chemical space 186 and to search large design spaces in a more systematic way to identify materials with desirable properties 77. These models work together to form an important foundation for the deep learning architectures that emerged later. While these classical and some early RL methods provided the foundational logic for data-driven research, they lacked the representation learning capabilities required to process multi-modal experimental data.
@@
The 2010s saw a shift away from feature engineering and toward representation learning. Large volumes of data and more readily available computing powered the development of deep learning 205, 132. The 2012 ImageNet competition winner, AlexNet194, showed that deep learning is effective as a general-purpose machine learning tool for real-world problems. In materials workflows, this time is less important as ”AI history” and more important as the time when learned representations started to take the place of manually created descriptors 394, 382. This shift enabled scalable screening, surrogate modeling, and downstream decision-making in materials discovery pipelines 162, 58.
After discriminative models proved useful and matured, the focus has been shifted to generative modeling. Generative Adversarial Networks (GANs) 133 made inverse design possible by training generator and discriminator networks against each other. Variational Autoencoders (VAEs) 190 had efficient inference and learning in probabilistic generative models with continuous latent variables.
Recurrent architectures such as Long Short-Term Memory (LSTM) 84 and Gated Recurrent Units (GRU) 75 have made temporal modeling for reaction prediction, synthesis pathway optimization, and material degradation forecasting possible. Gradient-based optimizers, such as Adam and SGD with momentum, together with regularization techniques like dropout 334 and batch normalization 157, helped make training more stable and made it easier to build deep models more reliably. Tools like TensorFlow 2 and PyTorch 275 played an important role in simplifying model development. Especially their support for multi-GPU training further accelerated progress and made it feasible to work with large databases of materials 176. As datasets continued to expand and recurrent architectures struggled with long-range dependencies, the community has tried to explore models better suited to handle global context. This shift has led to the rise of the attention-based Transformer era.
@@ 2.2 Transformer and Foundation Models
The self-attention mechanism 365 reshaped the sequence modeling tasks ever since it was proposed in the Transformer framework. This made it possible to perform parallel computation and model long-range dependencies, which are important for understanding chemical SMILES strings 386 and scientific literature. The multi-head attention 365 method lets models look at many representation subspaces at the same time, which lets them see how different representation subspaces interact in complex ways.
BERT (Bidirectional Encoder Representations from Transformers) 95 made bidirectional pre-training with masked language modeling and next sentence prediction in late 2018. This enabled models to capture the complex, context-dependent relationships that characterize scientific writing. GPT-3 56 demonstrated remarkable few-shot learning abilities with 175 billion parameters. Its performance suggested that scale together with data, can give models surprising flexibility, even without heavy fine-tuning.
@@
The term ”foundation models” was first used in 51 to describe large models trained on a variety of data, typically using self-supervision at scale, that can be applied to a wide range of tasks. This is because attention-based architectures made it possible for these models to be built. These models represent a major shift in AI: their unprecedented size leads to new abilities in language, vision, reasoning, and human interaction, but they also increase the risks of homogenization, where flaws or biases in the foundation model spread to all adapted downstream systems. Task-specialized Large Language Models (LLMs) trained or adapted on curated corpora (e.g., code 70, 257, math 209, 83, legal text 63) exemplify the dual nature of scaling: enhanced domain proficiency frequently accompanies hallucination 163, calibration, and faithfulness 213, alongside difficulties in controllability and verifiability 101.
@@
There are now foundation models that are made for a specific field that scientists can use. Domain-adapted scientific foundation models tailor general priors to expert corpora and evaluation regimes. SciBERT 37 modified BERT for scientific literature by training it on 1.14 million papers from Semantic Scholar. By improving skills like named entity recognition, citation prediction, and relation extraction, it made biomedical and computer science texts easier to understand. AlphaFold 174 made a big step forward in predicting protein structures with attention mechanisms for evolutionary data and physical constraints. This showed how transformers can be used for complex scientific problems, motivating analogous designs in other scientific domains (including materials science).
Large Language Models and Reasoning Capabilities
@@
Recent research has focused on reasoning, knowledge grounding, and tool-enhanced inference, alongside model scale. Chain-of-thought prompting 384 adds the scientific reasoning that is essential for material discovery. Memory-augmented architectures 52 support the retrieval of external knowledge, hence enhancing access to specialist materials science literature and databases during inference. Scaling laws 180 show that performance gets better when the number of parameters and data increases, following power-law relationships up to the limits of the regime studied. This makes materials research very challenging because high-quality datasets are far less abundant than in many other domains.
@@
In response to data scarcity, the community has explored strategies including the use of synthetic data such as OC20 66, pretraining on large unlabeled materials databases such as the Materials Project 160, and transfer learning from related scientific domains, including chemical datasets such as QM9 294. The inherent tension between data requirements and availability constrains foundation model applications in materials discovery to subdomains with sufficient training instances. The improvement in the reasoning ability of LLMs enhanced multiple fields in materials science research 407, such as enhanced reasoning, which can help molecular structure design 187 and material design 385.
@@
Cross-modal attention mechanisms have been used in multimodal foundation models 211, 325 to integrate information from text, images, and structured data, which is also useful for materials science, when workflows require fusing heterogeneous evidence (e.g., literature, microscopy, and structure representations) into a single decision context. Vision-language models (VLMs) like CLIP 292 and BLIP 210 use contrastive learning to enhance vision and language representation alignment and thereby provide retrieval-ready joint embeddings that can be repurposed with limited supervision. This ability has enabled tasks such as zero-shot classification and text-image retrieval for the materials characterization study (e.g., rapid triage, annotation, and cross-referencing of characterization images with textual descriptors), but reliability typically depends on domain-specific calibration and evaluation. Flamingo 13, KOSMOS-2 278, and VILA 219 are working on enhancing the video and 3D understanding abilities of the current multimodal framework, by extending multimodal context length and enabling richer image–text interleaving; however, robust 3D understanding for scientific data streams remains an open gap for end-to-end discovery settings.
In addition to general-purpose multimodal models, there are also specialized multimodal LLMs that have been adapted for materials science. These include MatterChat 349, MicroscopyGPT 80, MolVision 7, and PolyLLMem 417. These models combine atomic structure data, text, and images to make it easier for people and AI to work jointly on tasks such as discovering new materials and predicting their properties. However, the impact of their discovery depends on how well the system integrates aspects like tool use, uncertainty estimation, and closed-loop validation, not just the model’s ability in isolation.

### v2

@@ Abstract
The convergence of artificial intelligence and materials science presents a transformative opportunity, but achieving true acceleration in discovery requires moving beyond task-isolated, fine-tuned models toward agentic systems that plan, act, and learn across the full discovery loop. This survey advances a unique pipeline-centric view that spans from corpus curation and pre-training, through domain adaptation and instruction tuning, to goal-conditioned agents interfacing with simulation and experimental platforms. Unlike prior reviews, we treat the entire process as an end-to-end system to be optimized for tangible discovery outcomes rather than proxy benchmarks. This perspective allows us to trace how upstream design choices—such as data curation and training objectives—can be aligned with downstream experimental success through effective credit assignment.
To bridge communities and establish a shared frame of reference, we first present an integrated lens that aligns terminology, evaluation, and workflow stages across AI and materials science. We then analyze the field through two focused lenses: From the AI perspective, the survey details Large Language Model(LLM) strengths in pattern recognition, predictive analytics, and natural language processing for literature mining, materials characterization, and property prediction; from the materials science perspective, it highlights applications in materials design, process optimization, and the acceleration of computational workflows via integration with external tools (e.g., density functional theory (DFT), robotic labs). Finally, we contrast passive, reactive approaches with agentic design, cataloging current contributions while motivating systems that pursue long-horizon goals with autonomy, memory, and tool use. This survey charts a practical roadmap towards autonomous, safety‑aware LLM agents aimed at discovering novel and useful materials.
@@ 1 Introduction
Figure 1: An overview of the key sections and an illustrative end-to-end pipeline encompasses key elements like general pre-training tasks & data, foundation language models, domain-specific tasks & data, materials-oriented model adaptation, goal-driven Large Language Model (LLM) agents, and iterative open-ended materials experimentation. The colored arrows show how feedback signals from open-ended experiments are routed to the agentic LLM, the materials science-tuned LLM, the pre-trained LLM, and their corresponding task modules. Different colors of arrows indicate the sources of the feedbacks. For example, if general-purpose pre-training data does not contribute positively to the ultimate goal of novel and useful materials discovery, adjustments should be made by revising the pre-training tasks and corpus or fine-tuning the pre-trained LLM accordingly to better align model knowledge with the ultimate objective of discovering novel materials. Dashed arrows denote the forward information flows. Note that humans are not only responsible for monitoring the open-ended experiments, but also co-improve with the agentic LLM.
@@
Materials science faces the primary challenges of integrating insights from an expanding body of literature and ultimately expediting the autonomous discovery of novel functional materials. To solve these problems, we need to combine knowledge from different types of data, different scales, and different experimental settings. Conventional uses of machine learning in materials science, such as the initial use of LLMs, have mostly followed a static model: models are trained on a set of curated datasets to do specific tasks, like predicting properties or extracting entities W. Zhang, Q. Wang, X. Kong, J. Xiong, S. Ni, D. Cao, B. Niu, M. Chen, Y. Li, R. Zhang, et al. (2024b), 337, J. Van Herck, M. V. Gil, K. M. Jablonka, A. Abrudan, A. S. Anker, M. Asgari, B. Blaiszik, A. Buffo, L. Choudhury, C. Corminboeuf, et al. (2025). Pre‑trained LLMs excel at text understanding, efficiently mining chemical notation, experimental protocols, and technical jargon from unstructured text, thereby improving knowledge extraction and hypothesis generation from the literature 168, 200. Emerging multimodal frameworks that integrate textual, graph‑based, and image data enrich molecular and materials representations for retrosynthesis planning, catalyst design, and structure–property understanding across scales 223, 180, 407. Instruction‑tuning or other post-training approaches further bridge the gap between human intuition and machine execution by following complex, user‑defined workflows in materials research; yet, they remain constrained when operating without closed‑loop feedback and long‑horizon objectives. Consequently, progress toward autonomous discovery remains limited not by a lack of powerful models, but by the absence of an integrated framework that connects these isolated capabilities to improved experimental outcomes.
@@
This disconnection between static training proxies and dynamic utility reflects a broader pathology in AI evaluation. According to METR’s RE‑Bench analysis 393, frontier models resorted to sophisticated exploits, for example, monkey‑patching the time function to spoof evaluation. Echoing the “reward is enough” hypothesis 325 in reinforcement learning (RL), which argues that maximizing a suitably defined scalar reward can, in principle, drive the emergence of general intelligence, we posit that a carefully designed, end‑to‑end discovery reward, i.e., grounded in the successful identification and validation of novel, useful materials, is in principle sufficient to shape the entire AI4MatSci pipeline toward our ultimate scientific objective. It is not merely high benchmark performance but the autonomous generation of novel and useful materials through a result‑oriented, end-to‑end system capable of robust inference, adaptive experimentation, and safe deployment. In classical deep learning, end‑to‑end refers to the fact that a single scalar loss defined at the final task output is back‑propagated through all differentiable components of the model and feature stacks, enabling joint optimization of representation learning, intermediate modules, and prediction heads without manually engineered interfaces. Unfortunately, the traditional reactive and phase-centric approach confines the model’s capability to operate to individual stages of the entire workflow, making it difficult to adapt across dynamic, multi-modal, and potentially lifelong workflows that are essential for achieving autonomous materials discovery. Furthermore, today’s LLMs lack long‑horizon reasoning and seamless integration of multi‑scale, multimodal data streams, which are indispensable for driving an end-to-end materials discovery pipeline in realistic settings. Our central thesis extends this notion of end‑to‑end beyond a single neural network to the full materials discovery pipeline in Fig. 1, treating pre‑training, domain adaptation, tool use, experiment planning, and autonomous labs not as fixed stages but as coupled, trainable components linked by credit assignment propagated directly from real-world discovery outcomes.
@@
Van et al. 365 Broad Broad No Task‑based taxonomy
Madika et al. 237 Broad No LLMs Yes Physics‑informed, and responsible AI
Pyzer-Knapp et al. 288 Materials discovery for crystalline and molecular materials LLMs Yes LLMs for materials science
Peivaste et al. 279 Broad Broad No Emphasis on data
Jiang et al. 167 Text‑centric Text‑centric Yes Literature mining
Lei et al. 208 Broad Broad No LLMs augmenting human researchers
@@
Through the pipeline-centric lens, as shown in Fig. 1, this survey advocates shifting AI4MS from static fine-tuning to goal-oriented agent workflows for materials discovery. The discussion is structured around three guiding questions: Q1 (Evolution of Agency): How have general-purpose machine learning models evolved from merely passive processors to agentic systems? Q2 (Limitations of Phase-Centric Approaches) Why are existing AI for materials science systems not enough to meet the needs of autonomous materials discovery, even though they score well on material-related benchmarks? Q3 (The End-to-End Gap) What capacities have been developed, and how can we bridge the gaps in capability, optimization, and governance between existing AI for materials science and an end-to-end autonomous materials discovery pipeline?
There exist surveys on relevant topics, but they differ from ours, which are summarized in Tab. 1. In contrast, we present a unified pipeline-centric perspective for review and discussion. Driven by recent enablers such as autonomous labs, agent frameworks, and foundation models, this survey charts a practical roadmap from foundation models to autonomous, safety-aware LLM agents for discovering novel and useful materials that satisfy real performance, cost, and reliability criteria. Targeting both AI researchers and materials scientists, we bridge these disciplines to show why a system-level paradigm shift is now both necessary and feasible.
Figure 2: Technology tree of AI4Material science research. With the emergence of LLMs and agents, research on materials science initially focused on domain specific task, primarily concentrating on seperate reactive tasks. Subsequent research has delved deeper, gradually integrating more with agentic systems for materials science.
@@ 2 Recent Progress of AI and LLMs
The evolution from traditional machine learning to LLMs has progressively recast AI from a computational tool into an intelligent research collaborator capable of reasoning, generating hypotheses, orchestrating simulations, and designing experiments 289. In this section, we review recent AI and LLM advances with prospective relevance to materials science—irrespective of current deployment—to address Q1 in Sec. 1: How have general-purpose machine learning models evolved from merely passive processors to agentic systems? Rather than a chronological history, we trace this evolution through four key stages that mirror the pipeline in Fig. 1: (1) predictive models trained on static datasets; (2) foundation models serving as programmable priors; (3) post-training methods for objective shaping and controllability; and (4) agentic systems enabling tool use and long-horizon rewards. As shown in Fig. 1, pre-trained foundation models generally initialize the agentic pipeline for materials science; consequently, pre-training choices shape downstream capabilities and should be iteratively adapted using RL signals from real‑world experiments.
@@ 2.1 Deep Learning Revolution
From the late 1990s through the early 2010s, most early work in machine learning and AI was based on feature engineering, establishing the first batch of data-driven tools for materials science. Over this period, these early methods were not just used in materials research; they quickly became important instruments for finding new materials and predicting their properties.
There are three main types of machine learning, which have been applied to materials science. Supervised learning techniques, such as support vector machines 86, random forests 56, and shallow neural networks 21, facilitated property prediction from labeled data, thereby creating quantitative structure-property relationships for expedited materials screening.  Ward et al. 385 introduced a general-purpose framework that combined a chemically diverse feature set with ensemble-based models like random forests. Their framework worked well on both crystalline and amorphous systems, showing how powerful systematic feature design can be in speeding up the discovery of new materials. Unsupervised learning techniques, such as clustering (e.g., K-Means 242, DBSCAN 184) and dimensionality reduction (e.g., PCA 3, t-SNE 234), aided materials discovery by enabling pattern recognition, phase identification, and feature extraction from high-dimensional data 399. Both supervised and unsupervised learning algorithms enabled the initial requirements of the discovery cycle, i.e., correlating structures to properties.
To meet the need for autonomous materials discovery, RL techniques (e.g., Q-Learning 340, Deep Q-Networks (DQN) 248, Proximal Policy Optimization (PPO) 314, Model Predictive Control (MPC) 121, and Actor–Critic methods 192) were introduced to interact with an environment and learn from the consequences of their actions by optimizing reward functions over sequences of decisions. These methods make it possible to explore chemical space 187 and to search large design spaces in a more systematic way to identify materials with desirable properties 78. These models work together to form an important foundation for the deep learning architectures that emerged later. While these classical and early RL methods provided the foundational logic for data-driven research, they lacked the representation learning capabilities required to process multi-modal experimental data.
@@
The 2010s saw a shift away from feature engineering and toward representation learning. The availability of large-scale datasets and increased computational resources powered the development of deep learning 206, 133. The 2012 ImageNet competition winner, AlexNet 195, helped establish deep learning as a general-purpose machine learning tool for real-world problems. In materials workflows, this period matters less as a milestone in ”AI history” and more as the point when learned representations started to replace manually designed descriptors 397, 385. This shift enabled scalable screening, surrogate modeling, and downstream decision-making in materials discovery pipelines 163, 59.
After discriminative models proved useful and matured, the focus shifted to generative modeling. Generative Adversarial Networks (GANs) 134 made inverse design possible by training generator and discriminator networks against each other. Variational Autoencoders (VAEs) 191 had efficient inference and learning in probabilistic generative models with continuous latent variables.
Recurrent architectures such as Long Short-Term Memory (LSTM) 85 and Gated Recurrent Units (GRU) 76 have made temporal modeling for reaction prediction, synthesis pathway optimization, and materials degradation forecasting possible. Gradient-based optimizers, such as Adam and SGD with momentum, together with regularization techniques like dropout 336 and batch normalization 158 improved training stability and reliability. Tools like TensorFlow 2 and PyTorch 277 played an important role in simplifying model development. Especially their support for multi-GPU training further accelerated progress and made it feasible to work with large databases of materials 177. As datasets continued to expand and recurrent architectures struggled with long-range dependencies, the community has explored models better suited to handle global context. This shift has led to the rise of the attention-based Transformer era.
@@ 2.2 Transformer and Foundation Models
The self-attention mechanism 368 has reshaped sequence modeling tasks since it was proposed in the Transformer framework. This made it possible to perform parallel computation and model long-range dependencies, which are important for understanding chemical SMILES strings 389 and scientific literature. The multi-head attention 368 method enables models to look at many representation subspaces simultaneously, capturing complex interactions across them.
BERT (Bidirectional Encoder Representations from Transformers) 96 introduced bidirectional pre-training with masked language modeling and next sentence prediction in late 2018. This enabled models to capture the complex, context-dependent relationships that characterize scientific writing. GPT-3 57 demonstrated remarkable few-shot learning abilities with 175 billion parameters. Its performance suggested that scale, together with data, can give models surprising flexibility, even without heavy fine-tuning.
@@
The term ”foundation models” was first used in 52 to describe large models trained on a variety of data, typically using self-supervision at scale, that can be applied to a wide range of tasks. This is because attention-based architectures made it possible for these models to be built. These models represent a major shift in AI: their unprecedented size leads to new abilities in language, vision, reasoning, and human interaction, but they also increase the risks of homogenization, where flaws or biases in the foundation models spread to all adapted downstream systems. Task-specialized LLMs trained or adapted on curated corpora (e.g., code 71, 259, math 210, 84, legal text 64) exemplify the dual nature of scaling: enhanced domain proficiency frequently accompanies hallucination 164 and challenges in calibration and faithfulness 214, alongside difficulties in controllability and verifiability 102.
@@
Field-specific foundation models have also emerged for scientific use. Domain-adapted scientific foundation models tailor general priors to expert corpora and evaluation regimes. SciBERT 38 modified BERT for scientific literature by training it on 1.14 million papers from Semantic Scholar. By improving skills like named entity recognition, citation prediction, and relation extraction, it made biomedical and computer science texts easier to understand. AlphaFold 175 marked a major advance in predicting protein structures with attention mechanisms for evolutionary data and physical constraints. This showed how transformers can be used for complex scientific problems, motivating analogous designs in other scientific domains (including materials science).
LLMs and Reasoning Capabilities
@@
Recent research has focused on reasoning, knowledge grounding, and tool-enhanced inference in addition to scaling. Chain-of-thought prompting 387 adds the scientific reasoning that is essential for materials discovery. Memory-augmented architectures 53 support the retrieval of external knowledge, hence enhancing access to specialist materials science literature and databases during inference. Scaling laws 181 show that performance improves when the number of parameters and data increases, following power-law relationships up to the limits of the regime studied. This poses a challenge for materials research because high-quality datasets are far less abundant than in many other domains.
@@
In response to data scarcity, the community has explored strategies including the use of synthetic data such as OC20 67, pre-training on large unlabeled materials databases such as the Materials Project 161, and transfer learning from related scientific domains, including chemical datasets such as QM9 296. The inherent tension between data requirements and availability constrains applications of foundation models in materials discovery to subdomains with sufficient training instances. The improvement in the reasoning ability of LLMs enhanced multiple fields in materials science research 410, such as enhanced reasoning, which can help molecular structure design 188 and materials design 388.
@@
Cross-modal attention mechanisms have been used in multimodal foundation models 212, 327 to integrate information from text, images, and structured data, which is also useful for materials science, when workflows require fusing heterogeneous evidence (e.g., literature, microscopy, and structure representations) into a single decision context. Vision-language models (VLMs) like CLIP 294 and BLIP 211 use contrastive learning to enhance vision and language representation alignment and thereby provide retrieval-ready joint embeddings that can be repurposed with limited supervision. This ability has enabled tasks such as zero-shot classification and text-image retrieval for the materials characterization study (e.g., rapid triage, annotation, and cross-referencing of characterization images with textual descriptors), but reliability typically depends on domain-specific calibration and evaluation. Flamingo 13, KOSMOS-2 280, and VILA 221 are working on enhancing the video and 3D understanding abilities of the current multimodal framework, by extending multimodal context length and enabling richer image–text interleaving; however, robust 3D understanding for scientific data streams remains an open gap for end-to-end discovery settings.
In addition to general-purpose multimodal models, there are also specialized multimodal LLMs that have been adapted for materials science. These include MatterChat 351, MicroscopyGPT 81, MolVision 7, and PolyLLMem 420. These models combine atomic structure data, text, and images to make it easier for people and AI to work jointly on tasks such as discovering new materials and predicting their properties. However, the impact on discovery depends on how well the system integrates aspects like tool use, uncertainty estimation, and closed-loop validation, not just the model’s ability in isolation.